OpenStudy (18jonea):

The square root of a negative number is an imaginary number. True False @robtobey is this false

OpenStudy (18jonea):

@Michele_Laino

OpenStudy (michele_laino):

it is true! for example, we can write: \[\Large \sqrt { - 4} = \sqrt 4 \cdot \sqrt { - 1} = 2i\]